what is £35 - 10%
what is £92 - 10%
what is add 4% to £8
12300 add 4%

Answer by Alan3354(69443)

what is £35 - 10%
10% - 0.10
= 35 - 0.10*35
= 0.9*35
= 31.5
=================
what is £92 - 10%
Same as above.
==================
what is add 4% to £8
4% = 0.04
= 8 + 0.04*8
= 1.04*8
----------------
12300 add 4%
Same as above.